OpenStudy (anonymous):

Calculate the efficiency of person who is able to do 15 J of work by putting in 125 J of energy. Round the answer to the ones place. Also include the correct symbol for the unit.

OpenStudy (espex):

So you recall that efficiency is \(\frac{output}{input}*100\), what did you get?

OpenStudy (anonymous):

Aw I don't divide by 100

OpenStudy (espex):

You don't divide, you multiply.

OpenStudy (anonymous):

So would it be 12 cause you divide 15/125*100

OpenStudy (espex):

12%, exactly.

OpenStudy (anonymous):

Aw ok I get it now it's not a decimal I'm looking for it's a percent

OpenStudy (espex):

Right, so you get 0.12 as a decimal and then you multiply by 100 to get your percentage.

OpenStudy (anonymous):

Alright I see know but I did that to the next problem which is 2400/3100*100 and I got 77.4193

OpenStudy (espex):

Okay, I do not know the question so I can only say that if your output was 2400 and the input was 3100, then 77.4% efficiency looks right.

OpenStudy (anonymous):

The question is Calculate the efficiency of a machine that is able to do 2400 J of work when 3100 J of energy is added to it. Round your answer to the ones place and include the appropriate symbol with your answer.

OpenStudy (espex):

Looks good to me.

OpenStudy (anonymous):

Yea same here now what about this one If a person is 80% efficient in lifting an object, how much work was done when the person input 150 J of energy? This answer should be rounded to the tens place. It also needs to include the appropriate unit with the answer

OpenStudy (espex):

This is going to be the same equation, you will just have to solve it for the output.

OpenStudy (espex):

\(efficiency = \frac{output}{input}\)

OpenStudy (espex):

Plug in what you know and solve for the unknown.

OpenStudy (espex):

That should have read: \(efficiency = \frac{output}{input}*100\)

OpenStudy (anonymous):

So would I divide 80 by 150

OpenStudy (espex):

You first divide by 100, so you take your 80% back to a decimal.

OpenStudy (espex):

Then you are left with \(.80 = \frac{output}{input}\)

OpenStudy (espex):

So how would you get your input to the other side?

OpenStudy (anonymous):

take the .80 and multiply

OpenStudy (espex):

Yes. That will give you the work (or output) of the person.

OpenStudy (anonymous):

And that'll be the answer

OpenStudy (espex):

Yep.

OpenStudy (espex):

What did you get?

OpenStudy (anonymous):

So it would be .80*125 right

OpenStudy (espex):

Where did you get 125?

OpenStudy (anonymous):

Wrong I meant150

OpenStudy (espex):

Yes, .8*150

OpenStudy (anonymous):

And that's 120

OpenStudy (espex):

Exactly, good job.

OpenStudy (anonymous):

So it's just 120 not 120%

OpenStudy (espex):

Your input is not in a percentage, so it would only be 120% if you multiplied it by 100.

OpenStudy (anonymous):

Alright is it the same for this one A person is 67% efficient. If they are able to do 75 J of useful work, what is their total work or energy input? Round your answer to the tens place. Be sure to include the appropriate unit in the answer.

OpenStudy (espex):

Make sure your units on the previous question is Joules. For this question you need to take note that they are giving you the "output".

OpenStudy (espex):

So take your same equation, \(efficiency = \frac{output}{input}*100\), fill in the parts you know and solve for the part you don't.

OpenStudy (anonymous):

So what's the symbol for joules

OpenStudy (espex):

J

OpenStudy (anonymous):

Right I knew that

OpenStudy (anonymous):

I did this wrong didn't I 67*100*75

OpenStudy (espex):

Efficiency = 67% input = ? output = 75 \(67\% = \frac{75}{input}*100\)

OpenStudy (espex):

What do you do first?

OpenStudy (anonymous):

You would take 75*100

OpenStudy (espex):

I would take the 100 over so that I only had to work with \(.67 = \frac{75}{input}\)

OpenStudy (anonymous):

Alright and so I would then .67*75 right

OpenStudy (espex):

Think about that, how would you remove the 75 from the other side? What would you be left with? You want to get the "input" alone.

OpenStudy (anonymous):

I would divide?

OpenStudy (anonymous):

Or no

OpenStudy (espex):

Sure, just look at what you would have though: \(\frac{.67}{75} = \frac{1}{input}\)

OpenStudy (espex):

What if you multiplied by the "input"?

OpenStudy (espex):

\(input * .67 = 75\)

OpenStudy (anonymous):

So what's the input?

OpenStudy (espex):

Assuming you took my suggestion, what would be the next step to solve for the input?

OpenStudy (anonymous):

I can't multiply the .67 by 75 and I can't divide either so I have no clue

OpenStudy (espex):

Why can you not divide out the .67 from both sides?

OpenStudy (anonymous):

It gives me 0.0089

OpenStudy (espex):

Really? Because that's not what I get.

OpenStudy (anonymous):

What you get

OpenStudy (espex):

\(input = \frac{75}{.67}\)

OpenStudy (anonymous):

But I don't know what the input is so I took .67/76 and got that answer above

OpenStudy (espex):

You are dividing .67 by 76, though I presume you meant 75, which is not what the problem is.

OpenStudy (espex):

It looks like you are doing this: \[\frac{.67}{75}=\frac{1}{input}\] which means that the number you got is the "inverse" of the input. So if you take your answer and invert it: \(\frac{1}{your ~answer}\) you will get the correct number.

OpenStudy (anonymous):

so 0.008933333333 is my answer

OpenStudy (espex):

Which needs to be inverted.

OpenStudy (anonymous):

I did didn't I?

OpenStudy (espex):

Not if you have 0.0089333

OpenStudy (espex):

Slow down a bit and look at this. You had \[.67=\frac{75}{input}\] and you want to isolate input. Now the first logical step is to get it out of the denominator: \[input*~.67=\frac{75}{input}*input\rightarrow input*.67 = 75\] but you still want to isolate "input" so the second logical step is to multiply \(\frac{1}{.67}\)(or divde .67) to both sides\[\frac{1}{.67}*input*.67=75*\frac{1}{.67}\rightarrow input = \frac{75}{.67}\]

OpenStudy (espex):

So now you know that your input is equal to \(\frac{75}{.67}\)

OpenStudy (anonymous):

Right

OpenStudy (espex):

And what is 75 divided by .67?

OpenStudy (anonymous):

111.940

OpenStudy (espex):

YES! Very good. :D

OpenStudy (anonymous):

So what's next

OpenStudy (espex):

You're done.

OpenStudy (anonymous):

Aw for real that's it

OpenStudy (espex):

Yep, that's it.

OpenStudy (anonymous):

thanks again

OpenStudy (espex):

You're welcome.